A bullet of mass m moving with a horizontal velocity u strikes a stationary wooden block of mass M suspended by a string of length L=50 cm.The bullet emerges out the block with the speed u/4.If M=6m, the minimum value of u so that the block can complete the vertical circle?


1
Expert's answer
2021-11-11T11:59:56-0500

The speed to go round the circle:


"v=\sqrt {gL}."

Apply conservation of momentum:


"mu=m\frac u4+Mv,\\\space\\
u=\frac{4M}{3m}v=\frac{4M}{3m}\sqrt {gL}=17.7\text{ m/s}."